OpenStudy (anonymous):

If the varsity cheerleaders replace 8 uniforms and 6 poms, the cost is $378. If they replace 6 uniforms and 9 poms, the cost is $333. What is the cost of each uniform and each pom?

OpenStudy (anonymous):

let cost of 1 uniform= $x cost of 1 pom= $ y Then 8x+6y=378 ...(1) 6x+9y=333 ...(2) or 2x+3y=111 ...(3) 8x+12y=444 ...(4) (4) - (1) gives 6y=444-378=66 y=66/6=11 calculate x also.